∫ (π‘₯2Β²+ 5)Β³2π‘₯𝑑π‘₯



1
Expert's answer
2022-06-01T02:57:49-0400
"\int(x^2+5)^32xdx"

"u=x^2+5, du=2xdx"



"\int(x^2+5)^32xdx=\int u^3du=\dfrac{u^4}{4}+C""=\dfrac{(x^2+5)^4}{4}+C"